# The Magnetic Filling in MAD Simulations and its Impact on the Jet in M87

### v1

@@ 2.5 Flux Eruptions
We average the data over a time span, in which a MAD-state occurs in each simulation from a subset of three simulations: K.80, K.90 and K.99. In Fig. 7 we emphasize the mass accretion rate and the magnetic flux over the time spans under investigation for each simulation with a solid line. The time spans of interest in this study were chosen so that the magnetic fluxes through the event horizon have a similar profile, especially for the maximum and minimum amplitudes to be comparable, and the total time span to be $\Delta t=600\,\mathrm{M}$. These choices affect the comparability of the different models, especially when interpreting averaged plasma parameters and ray-traced images of azimuthally asymmetric structures, created by the cyclic MAD-states. The notion of better comparability in the indicated time intervals in Fig. 7 are based on the fact that the simulations undergo a very similar increase and decrease in accreted magnetic field, that at the peak of $\phi_{\mathrm{BH}}$ causes the arrestation of the disk. This arises from an advectively driven phase of accumulation of magnetic field in the vicinity of the BH, that then overreaches the threshold magnetic pressure to be accreted and instead outward diffusion of the accreted plasma takes over (Jacquemin-Ide et al. 2025). This is observed to occur in each model at very similar evolution states of the simulations (cf. Fig. 7). The build up of magnetic flux in the beginning of each interval is very similar, as every simulation accretes non-arrested for a very similar time span, as is given for the duration of the arrested states of the disks in all three simulations. The MAD-state (or magnetic flux eruption) is indicated by the decrease in the mass accretion rate (cf. Fig. 7) after both magnetic fluxes and mass accretion rates reach its peak in every simulation. We note, that these time intervals are the only ones within the whole simulations, where we see such strong similarities in the magnetic flux profiles, especially at those similar and importantly late evolution times ($\sim 17000\;\mathrm{M}$) in the simulations. Since the electromagnetic contribution of the jet power depends quadratically on $\phi_{\mathrm{BH}}$ (Tchekhovskoy et al. (2011)), the EM energy flux contributions to the jet can be ruled out to be a result of strongly differing $\phi_{\mathrm{BH}}$ over the three simulations.
Figure 7: Interval of investigation in order to compare the jet morphologies of all simulations. The magnetic flux $\phi_{\mathrm{BH}}$ describes a similar curve in height amplitude and width in time intervals for all simulations (indicated by solid lines). The dashed lines solely indicate time intervals that are not of interest for the discussions of the jet morphology.
@@
Table 2: Time intervals for the respective simulations that are used for most of the GRHMD analysis and completely used for the GRRT (see Fig. 7).
Figure 8: 2D-slices of the equatorial plane showing plasma-$\beta$, $\sigma$ and $T_{\mathrm{e}}$ of the peak MAD-state of the simulations K.80, K.90 and K.99. Note the extended, arrow-like tail of the flux tube in K.99 that is twice the length compared to K.90.
Figure 9: Meridional 2D slices of the peak MAD-state showing plasma-$\beta$, $\sigma$ and $T_{\mathrm{e}}$ of the simulations K.80, K.90 and K.99 (equal times as in Fig. 8 for each simulation, respectively). We show the $180^{\circ}$ azimuthally rotated cross-section of K.99, for comparison of the MAD-states.
We want to characterize the magnetic flux eruptions or MAD-states, since their impact on the morphology are imperative to consider in the subsequent analysis. We show respectively simultanioues snapshots of the meridional and equatorial planes in the GRMHD simulations, where the flux eruptions are at their respective peak in Fig. 8 and Fig. 15. We note, that the comparison of the snapshots lie outside the time intervals of comparison Tab. 2, nevertheless the MAD-states impact the simulations inside these intervals already and we show the peak MAD-states to emphasize these effects. We want to demonstrate the maximal impact, that the flux eruptions can have on the plasma parameters and on the radiative signatures. The timescales from the onset of these magnetically arrested states to their peak flux eruptions and finally termination of the arrested disk do not differ significantly across the three models. This in turn aids the comparability of the time-intervals (compare Tab. 2) chosen for time-dependent and -averaged quantities. We want to emphasize that although the amplitude of magnetic flux that is built up and released is very similar across all simulations (cf. Fig. 7) the flux tubes (cf. Porth et al. (2019)) do vary strongly in terms of their electron temperature, magnetization and kinetic to magnetic pressure ratio (see Fig. 8 and Fig. 15). In the simulation K.80 these magnetic eruptions occur two-sided in this instance, whereas both K.90 and K.99 are completely azimuthally asymmetric (cf. Fig. 8), and the one-sided flux eruptions are much more stable in the latter simulations. K.99 shows a significantly extended flux tube that is sufficiently hot to imprint clear signatures of thermal synchrotron radiation in the ray-tracings (see Sec. 3 & Fig. 21). The jet structure is also highly dependent on the MAD-states as these flux tubes are able to contribute to the jet morphology, significantly (see Fig. 15; K.99 is again rotated azimuthally by $180^{\circ}$, to compare the MAD-states), and especially introduce high degrees of azimuthal asymmetries, which is important to consider when we investigate time-averaged quantities and morphologies in the following sections.
@@ 2.6 Comparison of the Jet Structure
Figure 10: Upper row: $\sigma=1$ contours in xzxz-cross-section corresponding to K.80 (green lines) K.90 (blue lines) and K.99 (red lines) and the $1\sigma$ standard deviation in the $x$-directions over the time intervals listed in Tab. 2. Each column presents poloidal cross-sections at azimuthal angles in $45^{\circ}$ increments. Bottom row: Bernoulli parameter $hu_{t}=1.02$ contours, where we otherwise show the exact analogues for it as in the $\sigma=1$ contours.
@@
We investigate the jet structure and its variability by means of the magnetization $\sigma$ and the Bernoulli parameter $hu_{t}$. The contour at $\sigma=1\equiv\sigma_{\mathrm{cut}}$ defines our cut-off of the jet spine in the ray-tracings shown in Sec. 3 and can be interpreted as our inner boundary of the jet sheath. The Bernoulli parameter at $hu_{t}>1.02$ separates the outflowing plasma in the jet from inflowing or gravitationally bound plasma in the disk and wind, so that we define the $hu_{t}=1.02$ contour as the outer boundary of the jet sheath. In Fig. 10 we show meridional slices ($x$-$z$-plane) in $45^{\circ}$ increments in the azimuthal angle $\phi$ of the $\sigma_{\mathrm{cut}}=1$ contour (top row) and the Bernoulli parameter contour at $hu_{t}=1.02$ (bottom row), that are averaged in $x$-direction over the respective time intervals listed in Tab. 2 for each simulation. Moreover, to emphasize the structural variability we show the standard deviations in $x$-direction, that are indicated by the colored areas. Note that, by $x$-directions we denote every axis that is perpendicular to the BH spin vector (in our simulations aligned with the $z$-axis). We explicitly do not average over phi, since the MAD states that are undergone in the chosen time intervals (cf. Fig. 7) introduce strong azimuthal asymmetries that would otherwise smear out.
@@
In Fig. 10, the runs K.80 and K.90 trace similar jet boundaries, whereas K.99 manifests a consistently wider jet for $\phi=\{0^{\circ},45^{\circ}\}$ and a narrower jet for $\phi=\{90^{\circ},135^{\circ}\}$. This hints at a rather elliptical paraboloid jet geometry in the K.99 run, likely induced by the stronger MAD-state that causes the azimuthal axis asymmetry. We note that we have rotated, K.99 azimuthally by $180^{\circ}$ such that the flux eruptions of the simulations match in azimuthal angle. Interestingly, at $\phi=135^{\circ}$ all runs consistently show the strongest axis asymmetries, which coincides roughly with the foot point of the flux eruption (compare 8, where the x-Axis is at $\phi=0^{\circ}/180^{\circ}$).
@@
The $1\sigma$-band in Fig. 10 shows consistent structural properties as the Bernoulli contour. The sigma contour is overall significantly narrower downstream the jet and less variable for K.99, especially at distances $\sim 100\mathrm{M}-200\mathrm{M}$ from the center, compared to K.80 and K.90.
3 General Relativistic Radiative Transfer
@@
In order to investigate the radiative signatures of the 3D GRMHD simulations we calculate the polarized synchrotron emission by means of a general relativistic radiative transfer (GRRT) post-processing step. We therefore employ IPOLE (Mościbrodzka & Gammie (2017)): a semi-analytic code for relativistic polarized radiative transfer.
@@
We assume a model for the electron Distribution Function (eDF), since our GRMHD simulations do neither evolve the electrons simultaneously nor separately, rather they are assumed to tightly follow the evolution of the simulated plasma. The model chosen for this study combines the thermal Maxwell-Jüttner distribution
$\frac{1}{n_{e}}\frac{dn_{e}}{d\Gamma_{\mathrm{e}}}=\frac{\Gamma_{\mathrm{e}}^{2}\sqrt{1-1/\Gamma_{\mathrm{e}}^{2}}}{\Theta_{e}K_{2}(1/\Theta)}\\$ (24)
@@
and the non-thermal $\kappa$-eDF (see Pandya et al. (2016)), that smoothly converges to the Maxwell-Jüttner distribution for low $\Gamma_{\mathrm{e}}$
@@
with the normalization factor $N$, the dimensionless electron temperature $\Theta_{e}=k_{b}T_{e}/m_{e}c^{2}$, the modified Bessel-function of the second kind $K_{2}$, the power-law index $\kappa$ and the width parameter $w$.
The thermal eDF is the relativistic generalization of the thermal Maxwell distribution and the $\kappa$-eDF is inspired by observations of the solar wind and by results of collisionless plasma simulations (e.g. Kunz et al. (2015)) and converges to the thermal distribution for $\kappa\rightarrow\infty$. Both distributions need an assignment of the electron temperature $T_{e}$, which we define as
@@
where $k_{\mathrm{B}}$, $m_{\mathrm{p}}$ and $m_{\mathrm{e}}$ are respectively the Boltzmann constant and the proton and electron masses. $\rho$ and $p$ are the rest-mass density and pressure from the GRMHD simulation, respectively. The code temperature $T_{\mathrm{code}}=\nicefrac{{p}}{{\rho}}$ is thus defined by plasma parameters from the GRMHD code as well and describes solely the ion temperature in code units. The ratio of proton temperature $T_{\mathrm{p}}$ and electron temperature $T_{\mathrm{e}}$, $R$, is defined as
@@
Figure 11: Time-averaged $86$ GHz ray-tracings of the M87 jet. Each simulation is averaged over the time intervals corresponding to each simulation that are listed in Tab. 2. Stokes-I is super-imposed with a line integral convolution of the EVPA vector field by scaling the opacity as a function of the linear polarization fraction and rescaling the image by rescaling the maximum of the convolved image $\max{[f(I,\,m,\,\mathrm{EVPA}_{\mathrm{vector-field}})]}$ to $\max{(\mathrm{I})}$.
@@
The non-thermal contributions to the electron energies are thought to be originated in magnetic reconnection events, in which current sheets accelerate the charged particles. These processes cannot be resolved in global GRMHD simulations, therefore, the $\kappa$-parameter depends on the magnetization and plasma-$\beta$ when employing a PIC simulation based sub-grid model from Ball et al. (2018):
@@ 3.1 M87 Ray-Tracing Setup
For the 86 GHz Ray-Tracings, we fixed the parameters of all runs to the mass of the BH of $M_{\mathrm{BH}}=6.2\times 10^{9}\,\mathrm{M_{\odot}}$ (cf. Gebhardt et al. 2011; EHT Collaboration et al. 2019a, b, c), the distance to the BH to $D=16.9\,\mathrm{Mpc}$ (cf. Blakeslee et al. 2009; Cantiello et al. 2018; EHT Collaboration et al. 2019b, c) and the viewing angle $i=163^{\circ}$ (as spanned by the BH spin axis and line of sight) of the BH and jet in M87. The parameters for the electron temperature model are set to $R_{\mathrm{high}}=160$ and $R_{\mathrm{low}}=1$ and for the $\kappa$-eDF in Eq. (25), we choose a non-thermal emission fraction of $\varepsilon=0.5$ and a injection radius of non-thermal particles at $r_{\mathrm{inj}}=10\,\mathrm{M}$. We note that the numerical approximations of the emission and absorption coefficients are only valid in the range $3.1\geq\kappa\leq 8$, where solely the $\kappa$-eDF is applied, and the Maxwell-Jüttner eDF otherwise.
@@
The time-averaged flux of each ray-traced simulation is adjusted such that the flux at horizon scale at 230 GHz matches the observed $0.9\,\mathrm{Jy}$ by the EHT Collaboration et al. (2019a) by means of optimizing the mass-accretion rates in the GRMHD simulations.
Figure 12: Top row: 86 GHz light curve for simulations K.80 (green) K.90 (blue) and K.99 (red) in the time intervals shown in Tab. 2 but shifted by an offset for reasons of better comparability. Bottom row: Linear Polarization fraction with the same color-code as before.

### v2

@@ 2.5 Flux eruptions
We averaged the data over a time span in which a MAD state occurs in each simulation from a subset of three simulations: K.80, K.90, and K.99. In Fig. 7, we use a solid line to emphasize the mass-accretion rate and the magnetic flux over the time spans under investigation for each simulation. The time spans of interest in this study were chosen so that the magnetic fluxes through the event horizon have a similar profile, especially for the maximum and minimum amplitudes to be comparable and the total time span to be $\Delta t=600\,\mathrm{M}$. These choices affected the comparability of the different models, especially when interpreting averaged plasma parameters and ray-traced images of azimuthally asymmetric structures, created by the cyclic MAD states. The notion of better comparability in the indicated time intervals in Fig. 7 is based on the fact that the simulations undergo very similar increases and decreases in accreted magnetic field, which, at the peak of $\phi_{\mathrm{BH}}$, causes the arrestation of the disk. This arises from an advectively driven phase of accumulation of magnetic field in the vicinity of the BH, which then overreaches the threshold magnetic pressure to be accreted, and outward diffusion of the accreted plasma takes over (Jacquemin-Ide et al. 2025). This is observed to occur in each model at very similar evolution states of the simulations (cf. Fig. 7). The build up of magnetic flux in the beginning of each interval is very similar, as every simulation accretes in a non-arrested way over a very similar time span, as is given for the duration of the arrested states of the disks in all three simulations. The MAD state (or magnetic-flux eruption) is indicated by the decrease in the mass-accretion rate (cf. Fig. 7) after both magnetic fluxes and mass-accretion rates reach their peak in every simulation. We note that these time intervals are the only ones within the whole simulations, where we see such strong similarities in the magnetic flux profiles, especially at those similar and importantly late evolution times ($\sim 17000\;\mathrm{M}$) in the simulations. Since the electromagnetic contribution of the jet power depends quadratically on $\phi_{\mathrm{BH}}$ (Tchekhovskoy et al. (2011)), the EM energy flux contributions to the jet can be ruled out from being a result of strongly differing $\phi_{\mathrm{BH}}$ over the three simulations.
Figure 7: Interval of investigation in order to compare the jet morphologies of all simulations. The magnetic flux, $\phi_{\mathrm{BH}}$, describes a similar curve in height amplitude and width in time intervals for all simulations (indicated by solid lines). The dashed lines solely indicate time intervals that are not of interest for the discussions of the jet morphology.
@@
Table 2: Time intervals for the respective simulations that were used for parts of the GRHMD analysis and only used for the GRRT.
Figure 8: 2D slices of the equatorial plane showing plasma $\beta$, $\sigma,$ and $T_{\mathrm{e}}$ of the peak MAD state of the simulations K.80, K.90, and K.99. Note the extended, arrow-like tail of the flux tube in K.99 that is twice the length compared to that of K.90.
Figure 9: Meridional 2D slices of peak MAD state showing plasma $\beta$, $\sigma,$ and $T_{\mathrm{e}}$ of the simulations K.80, K.90, and K.99 (equal times to those in Fig. 8 for each simulation, respectively). We show the $180^{\circ}$ azimuthally rotated cross-section of K.99, for comparison of the MAD states.
We wanted to characterize the magnetic flux eruptions or MAD states, since their impacts on the morphology are essential for consideration in the subsequent analysis. We show, respectively, simultaneous snapshots of the meridional and equatorial planes in the GRMHD simulations, where the flux eruptions are at their respective peaks in Fig. 8 and Fig. 9. We note that the comparison of the snapshots lies outside the time intervals of comparison in Tab. 2; nevertheless, the MAD states already impact the simulations inside these intervals, and we show the peak MAD states to emphasize these effects. We want to demonstrate the maximal impact, that the flux eruptions can have on the plasma parameters and on the radiative signatures. The timescales from the onset of these magnetically arrested states to their peak flux eruptions and, finally, termination of the arrested disk do not differ significantly across the three models. This in turn aids the comparability of the time intervals (compare Tab. 2) chosen for time-dependent and time-averaged quantities. We want to emphasize that although the amplitude of magnetic flux that is built up and released is very similar across all simulations (cf. Fig. 7), the flux tubes (cf. Porth et al. (2019)) do vary strongly in terms of their electron temperature, magnetization, and kinetic-to-magnetic pressure ratio (see Fig. 8 and Fig. 9). In the simulation K.80, these magnetic eruptions occur two-sided in this instance, whereas both K.90 and K.99 are completely azimuthally asymmetric (cf. Fig. 8), and the one-sided flux eruptions are much more stable in the latter simulations. K.99 shows a significantly extended flux tube that is hot enough to imprint clear signatures of thermal synchrotron radiation in the ray tracings (see Sect. 3 and Fig. 21). The jet structure is also highly dependent on the MAD states as these flux tubes are able to contribute to the jet morphology significantly (see Fig. 9; K.99 is again rotated azimuthally by $180^{\circ}$ to compare the MAD states), and, especially, they introduce high degrees of azimuthal asymmetries, which is important to consider when we investigate time-averaged quantities and morphologies in the following sections.
@@ 2.6 Comparison of the jet structure
Figure 10: Upper row: $\sigma=1$ contours in xzxz-cross-section corresponding to K.80 (green lines), K.90 (blue lines), and K.99 (red lines), and the $1\sigma$ standard deviation in the $x$-directions over the time intervals listed in Tab. 2. Each column presents poloidal cross-sections at azimuthal angles in $45^{\circ}$ increments. Bottom row: Bernoulli parameter $hu_{t}=1.02$ contours, where we otherwise show the exact analogs for it as in the $\sigma=1$ contours.
@@
We investigated the jet structure and its variability by means of the magnetization, $\sigma,$ and the Bernoulli parameter, $hu_{t}$. The contour at $\sigma=1\equiv\sigma_{\mathrm{cut}}$ defines our cutoff of the jet spine in the ray tracings shown in Sect. 3 and can be interpreted as our inner boundary of the jet sheath. The Bernoulli parameter at $hu_{t}>1.02$ separates the outflowing plasma in the jet from inflowing or gravitationally bound plasma in the disk and wind, so we defined the $hu_{t}=1.02$ contour as the outer boundary of the jet sheath. In Fig. 10, we show meridional slices ($x$-$z$-plane) in $45^{\circ}$ increments in the azimuthal angle, $\phi,$ of the $\sigma_{\mathrm{cut}}=1$ contour (top row) and the Bernoulli parameter contour at $hu_{t}=1.02$ (bottom row), which are averaged in the $x$-direction over the respective time intervals listed in Tab. 2 for each simulation. Moreover, to emphasize the structural variability we show the standard deviations in the $x$-direction, indicated by the colored areas. Note that by $x$-directions we mean every axis that is perpendicular to the BH spin vector (in our simulations aligned with the $z$-axis). We explicitly did not average over phi, since the MAD states that are undergone in the chosen time intervals (see Fig. 7) introduce strong azimuthal asymmetries that would otherwise smear out.
@@
In Fig. 10, the runs K.80 and K.90 trace similar jet boundaries, whereas K.99 manifests a consistently wider jet for $\phi=\{0^{\circ},45^{\circ}\}$ and a narrower jet for $\phi=\{90^{\circ},135^{\circ}\}$. This hints at a rather elliptical paraboloid jet geometry in the K.99 run, likely induced by the stronger MAD state that causes the azimuthal axis asymmetry. We note that we rotated K.99 azimuthally by $180^{\circ}$, such that the flux eruptions of the simulations match in azimuthal angles. Interestingly, at $\phi=135^{\circ}$ all runs consistently show the strongest axis asymmetries, which coincides roughly with the foot point of the flux eruption (compare to 8, where the x-axis is at $\phi=0^{\circ}/180^{\circ}$).
@@
The $1\sigma$ band in Fig. 10 shows consistent structural properties as the Bernoulli contour. The sigma contour is overall significantly narrower downstream of the jet and less variable for K.99, especially at distances of $\sim 100\mathrm{M}-200\mathrm{M}$ from the center, compared to K.80 and K.90.
3 General-relativistic radiative transfer
@@
In order to investigate the radiative signatures of the 3D GRMHD simulations, we calculated the polarized synchrotron emission by means of a general-relativistic radiative-transfer (GRRT) post-processing step. We therefore employed IPOLE (Mościbrodzka and Gammie (2017)): a semi-analytic code for relativistic polarized radiative transfer.
@@
We assumed a model for the electron distribution function (eDF), since our GRMHD simulations evolve the electrons neither simultaneously nor separately; rather, they are assumed to tightly follow the evolution of the simulated plasma. The model chosen for this study combines the thermal Maxwell-Jüttner distribution,
$\frac{1}{n_{e}}\frac{dn_{e}}{d\Gamma_{\mathrm{e}}}=\frac{\Gamma_{\mathrm{e}}^{2}\sqrt{1-1/\Gamma_{\mathrm{e}}^{2}}}{\Theta_{e}K_{2}(1/\Theta)}\\ ,$ (24)
@@
and the nonthermal $\kappa$-eDF (see Pandya et al. (2016)), which smoothly converges to the Maxwell–Jüttner distribution for low $\Gamma_{\mathrm{e}}$:
@@
with the normalization factor $N$; the dimensionless electron temperature $\Theta_{e}=k_{b}T_{e}/m_{e}c^{2}$; the modified Bessel function of the second kind, $K_{2}$; the power-law index $\kappa;$ and the width parameter $w$.
The thermal eDF is the relativistic generalization of the thermal Maxwell distribution, and the $\kappa$-eDF was inspired by observations of the solar wind and by results of collisionless plasma simulations (e.g. Kunz et al. (2015)) and converges to the thermal distribution for $\kappa\rightarrow\infty$. Both distributions need an assignment of the electron temperature $T_{e}$, which we define as
@@
where $k_{\mathrm{B}}$, $m_{\mathrm{p}}$, and $m_{\mathrm{e}}$ are, respectively, the Boltzmann constant and the proton and electron masses. $\rho$ and $p$ are the rest-mass density and pressure from the GRMHD simulation, respectively. The code temperature, $T_{\mathrm{code}}=\nicefrac{{p}}{{\rho}}$, is thus defined by plasma parameters from the GRMHD code as well, and it describes solely the ion temperature in code units. The ratio of proton temperature $T_{\mathrm{p}}$ and electron temperature $T_{\mathrm{e}}$, $R$, is defined as
@@
Figure 11: Time-averaged $86$ GHz ray tracings of the M87 jet. Each simulation is averaged over the time intervals corresponding to each simulation that are listed in Tab. 2. Stokes I is superimposed with a line integral convolution of the EVPA vector field by scaling the opacity as a function of the linear polarization fraction and rescaling the image by the maximum of the convolved image $\max{[f(I,\,m,\,\mathrm{EVPA}_{\mathrm{vector-field}})]}$ to $\max{(\mathrm{I})}$.
@@
The non-thermal contributions to the electron energies are thought to be originated in magnetic reconnection events, in which current sheets accelerate the charged particles. These processes cannot be resolved in global GRMHD simulations, therefore, the $\kappa$-parameter depends on the magnetization and plasma $\beta$ when employing a PIC simulation based sub-grid model from Ball et al. (2018):
@@ 3.1 M87 ray-tracing setup
For the 86 GHz Ray-Tracings, we fixed the parameters of all runs to the mass of the BH of $M_{\mathrm{BH}}=6.2\times 10^{9}\,\mathrm{M_{\odot}}$ (cf. Gebhardt et al. 2011; EHTC et al. 2019a, c, b), the distance to the BH to $D=16.9\,\mathrm{Mpc}$ (cf. Blakeslee et al. 2009; Cantiello et al. 2018; EHTC et al. 2019c, b) and the viewing angle $i=163^{\circ}$ (as spanned by the BH spin axis and line of sight) of the BH and jet in M87. The parameters for the electron temperature model are set to $R_{\mathrm{high}}=160$ and $R_{\mathrm{low}}=1$ and for the $\kappa$-eDF in Eq. (25), we choose a non-thermal emission fraction of $\varepsilon=0.5$ and a injection radius of non-thermal particles at $r_{\mathrm{inj}}=10\,\mathrm{M}$. We note that the numerical approximations of the emission and absorption coefficients are only valid in the range $3.1\geq\kappa\leq 8$, where solely the $\kappa$-eDF is applied, and the Maxwell-Jüttner eDF otherwise.
@@
The time-averaged flux of each ray-traced simulation is adjusted such that the flux at horizon scale at 230 GHz matches the observed $0.9\,\mathrm{Jy}$ by the EHTC et al. (2019a) by means of optimizing the mass-accretion rates in the GRMHD simulations.
Figure 12: Top row: 86 GHz light curve for simulations K.80 (green), K.90 (blue), and K.99 (red) in the time intervals shown in Tab. 2, but shifted by an offset for better comparability. Bottom row: Linear polarization fraction with the same color-code as before.